# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 3.2. Steenrod Operations
The Steenrod operations are cohomology operations widely used in algebraic topology. They can be sensibly defined on $S_{m}$ and restrict to operations on $S_{m}^{G_{m}}$. Thus applying a Steenrod operation to an invariant produces a potentially new invariant. The complete Steenrod operator $\mathcal{P}(t):S_{m}\to S_{m}[t]$ is the algebra homomorphism determined by $\mathcal{P}(t)(v)=v+v^{q}t$ for $v$ homogeneous of degree one. Since the map is linear in degree one, $\mathcal{P}(t)$ is well-defined. For $f$ homogeneous of degree $d$, the Steenrod operations $\mathcal{P}^{i}(f)$ are defined by
@@
Suppose $v,f\in S_{m}$ with $v$ homogeneous of degree one. Then $v$ divides $\mathcal{P}^{i}(vf)$ for all $i$.
@@
By definition, $v$ divides $\mathcal{P}^{j}(v)$. Therefore, using the Cartan identity, $v$ divides $\mathcal{P}^{i}(vf)$. ∎
It is an immediate consequence of the lemma that if $f$ is a product of linear forms, $f$ divides $\mathcal{P}^{i}(f)$.
@@
(a) $\mathcal{P}(t)(\xi_{0})=\xi_{0}+\xi_{1}t+\xi_{0}^{q}t^{2}$.
(b) $\mathcal{P}(t)(\xi_{1})=\xi_{1}+2\xi_{0}^{q}t+\xi_{2}t^{q}+\xi_{1}^{q}t^{q+1}$.
(c) For $i>1$, $\mathcal{P}(t)(\xi_{i})=\xi_{i}+\xi_{i-1}^{q}t+\xi_{i+1}t^{q^{i}}+\xi_{i}^{q}t^{q^{i}+1}$.
@@
(b) $\mathcal{P}^{j+1}(\xi_{0}b)=\xi_{1}b^{q}+\xi_{0}^{q}\mathcal{P}^{j-1}(b)$.
(c) $\mathcal{P}^{j+3}(\xi_{0}^{2}b)=2\xi_{0}^{q}\xi_{1}b^{q}+\xi_{0}^{2q}\mathcal{P}^{j-1}(b)$.
(d) $\mathcal{P}^{2q+j-1}(\xi_{0}^{q}b)=\xi_{0}^{q^{2}}\mathcal{P}^{j-1}b$.
@@ 3.7. New Invariants
Let $P_{m}$ denote the intersection of $G_{m}$ with the upper triangular unipotent subgroup of ${{\text{GL}_{{n}}(\mathbb{F}_{q})}}$. $P_{m}$ is a choice of Sylow $p$-subgroup for $G_{m}$. For $v\in S_{m}$ homogenous of degree one, define $N(v)$ to be the orbit product $v$ over $P_{m}$. Since $x_{1}\in S_{m}^{P_{m}}$, $N(x_{1})=x_{1}$. We will see in Section 5 that $\deg(N(x_{i}))=q^{i-1}$ and $\deg(N(y_{i}))=q^{n-i-1}$. Define
@@
We observe that each linear form that occurs as a factor of $u_{m}$ is isotropic under the induced symmetric bilinear form, $\beta^{*}$, on $V^{*}$. Furthermore every isotropic linear form is a scalar multiple of one of the factors of $u_{m}$.
Since $u_{m}$ is a product of linear forms, Lemma 3.4 implies that $u_{m}$ divides $\mathcal{P}^{j}(u_{m})$ in $S_{m}$. Define $e(i,m):=\sum_{j=1}^{i}q^{n-1-j}$ with $n=2m$ and, for $1\leq i\leq m$, define $d_{i,m}\in S_{m}$ by
@@
For convenience, define $d_{0,m}=1$. Observe that $e(i,m)=q^{n-2}+qe(i-1,m-1)$.
@@
$\psi(u_{m})/u_{m}$ is the minimal polynomial of $x_{1}$ over the field $\mathcal{F}(S_{m}^{G_{m}})$. In particular, $d_{i,m}\in S_{m}^{G_{m}}$ for $i=1,\ldots,m$.
@@
If we identify $G_{m-1}$ with the pointwise stabiliser of $\{y_{1},x_{1}\}$ in $G_{m}$, then the pointwise stabiliser of $x_{1}$ coincides with $G_{m-1}H$, where $H$ is the associated Hook group (see Section 5.1). From this we see that the size the orbit of $x_{1}$ is
$\frac{|{{\text{O}^{+}_{{n}}(\mathbb{F}_{q})}}|}{|{{\text{O}^{+}_{{n-2}}(\mathbb{F}_{q})}}||H|}=\frac{(q^{m}-1)(q^{2m-2})}{q^{m-1}-1}=(q^{m}-1)(q^{m-1}+1)=(q-1)\deg(u_{m})$
and this orbit consists of the non-zero scalar multiples of the linear factors of $u_{m}$. If $v$ is a linear factor of $u_{m}$, then the roots of $\psi(v)$ are the non-zero scalar multiples of $v$. From this, using the fact that $\psi(u_{m})/u_{m}$ is monic of degree $(q-1)\deg(u_{m})$, we conclude that $\psi(u_{m})/u_{m}=\prod\{t-x_{1}g\mid g\in G_{m}\}\in S_{m}^{G_{m}}$. Therefore, $\psi(u_{m})/u_{m}$ is the minimal polynomial of $x_{1}$ over $\mathcal{F}(S_{m}^{G_{m}})$ and, since the $d_{i,m}$ are coefficients of $\psi(u_{m})/u_{m}$ up to sign, we have $d_{i,m}\in S_{m}^{G_{m}}$. ∎
@@
$\operatorname{LT}(u_{m})=\prod_{j=1}^{m}y_{j}^{q^{n-j-1}}x_{i}^{q^{j-1}}$
and $\operatorname{LT}(d_{1,m})=y_{1}^{q^{n-1}-q^{n-2}}$.
@@
$u_{m}d_{1,m}=\mathcal{P}^{q^{n-2}}\left(N(y_{1})\right)x_{1}\prod_{i=2}^{m}N(y_{i})N(x_{i})+F$
where $F$ is smaller than $\operatorname{LT}(\mathcal{P}^{q^{n-2}}(N(y_{1}))=y_{1}^{q^{n-1}}$ in the lexicographic order. Dividing by $u_{m}$ gives $\operatorname{LT}(d_{1,m})=y_{1}^{q^{n-1}-q^{n-2}}$. ∎
Define :=$\mathcal{H}:=\{\xi_{0},\ldots,\xi_{m-1},d_{1,m},\ldots,d_{m,m}\}$.
@@
$\mathcal{V}(\xi_{0},\ldots,\xi_{m-1})=\bigcup_{g\in{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}g\mathcal{V}(x_{1},x_{2},\ldots,x_{m}).$
For $v\in\mathcal{V}(\xi_{0},\ldots,\xi_{m-1})$, choose $g\in{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ so that $gv\in\mathcal{V}(x_{1},x_{2},\ldots,x_{m})$. From Theorem 3.9, $d_{i,m}\in S_{m}^{G_{m}}$. Therefore $d_{i,m}(gv)=d_{i,m}(v)$. Hence to show $\mathcal{V}(\mathcal{H})=\{\underline{0}\}$, it is sufficient to show $\mathcal{V}(x_{1},\ldots,x_{m},d_{1,m},\ldots,d_{m,m})=\{\underline{0}\}$. To do this we consider $d_{i,m}$ modulo the ideal $I:=\langle x_{1},x_{2},\ldots,x_{m}\rangle$.
Define $\bar{d}_{i}\in\mathbb{F}_{q}[y_{1},\ldots,y_{m}]$ by $\bar{d}_{i}\equiv_{I}d_{i,m}$. We will show that $\{\bar{d}_{1},\ldots,\bar{d}_{m}\}$ is an hsop in $\mathbb{F}_{q}[y_{1},\ldots,y_{m}]$. Writing $d_{1,m}$ in terms of the generators of the Borel invariants (see Section 5.18), shows that $d_{1,m}$ is congruent to an element of the ring $\mathbb{F}_{q}[N(y_{1})^{q-1},\ldots,N(y_{m})^{q-1}]$. Define $W_{i}:={\rm Span}_{\mathbb{F}_{q}}\{y_{i},\ldots,y_{m}\}$ and observe that
$N(y_{i})\equiv_{I}\prod_{w\in W_{i+1}}(y_{i}+w)^{q^{m-1}}=\prod_{w\in W_{i+1}}(y_{i}^{q^{m-1}}+w^{q^{m-1}})$
(for $i=m$, $W_{m+1}=0$ and $N(y_{m})\equiv_{I}y_{m}^{q^{m-1}}$). Therefore $\bar{d}_{1}\in\mathbb{F}_{q}[y_{1}^{q^{m-1}},\ldots,y_{m}^{q^{m-1}}]$. Observe that taking $g$ to $g\oplus(g^{-1})^{T}$ gives an embedding of ${{\text{GL}_{{m}}(\mathbb{F}_{q})}}$ into ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$. Using this embedding and the fact that $d_{1,m}\in S_{m}^{G_{m}}$, we have
$\bar{d}_{1}\in\mathbb{F}_{q}[y_{1}^{q^{m-1}},\ldots,y_{m}^{q^{m-1}}]^{{{\text{GL}_{{m}}(\mathbb{F}_{q})}}}\ .$
By comparing lead terms (see Lemma 3.10), we see that
$\bar{d}_{1}=d_{1}(y_{1}^{q^{m-1}},\ldots,y_{m}^{q^{m-1}})=d_{1}(W_{1})^{q^{m-1}},$
where $d_{1}(W_{1})$ is the first Dickson invariant for $W_{1}$. For a definition and background material on the Dickson invariants see [24] or [12] or [5, §3.3]. We claim that
$d_{i,m}\equiv_{I}\pm d_{i}(W_{1})^{q^{m-1}}$
for all $i$. From this, it follows that $\{\bar{d}_{1},\ldots,\bar{d}_{m}\}$ is an hsop for $\mathbb{F}_{q}[y_{1},\ldots,y_{m}]$, as required.
To see that $d_{i,m}\equiv_{I}\pm d_{i}(W_{1})^{q^{m-1}}$, we look at $\psi(u_{m})/u_{m}$ modulo $I$:
$\psi(u_{m})/u_{m}=\prod\{t-x_{1}g\mid g\in G_{m}\}\equiv_{I}t^{q^{m}-1}\left(\prod_{w\in W_{1}\setminus\{0\}}(t-w)\right)^{q^{m-1}}.$
On the right-hand side of this equivalence, $d_{i}(W_{1})^{q^{m-1}}$ is the coefficient of
@@
from which we conclude $d_{i,m}\equiv_{I}\pm d_{i}(W_{1})^{q^{m-1}}$, where the parity is determined by $(-1)^{\ell}=(-1)^{i+1}$. ∎
@@
$\operatorname{LT}(d_{i,m})=\prod_{j=1}^{i}y_{j}^{q^{n-j}-q^{n-j-1}}.$
@@ 4. Formulating the Main Theorem
Define $R_{i}:=\mathbb{F}_{q}[T_{i},T_{i-1},\ldots,T_{0}]$ and let $\Phi_{i,m}:R_{i}\to S_{m}$ denote the algebra map defined by $\Phi_{i,m}(T_{j})=\xi_{j}$. If we grade $R_{i}$ by assigning $T_{j}$ degree $q^{j}+1$, then $\Phi_{i,m}$ preserves degree. Since $\{\xi_{0},\ldots,\xi_{2m-1}\}$ is algebraically independent in $S_{m}$, $\Phi_{i,m}$ is injective for $i<2m$.
@@
The kernel of $\Phi_{2m,m}$ is a principal ideal. Furthermore, the generator is of the form $aT_{2m}-b$ with $a,b\in R_{2m-1}\setminus\{0\}$.
@@
The field of fractions of $S_{m}^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}$ is $\mathbb{F}_{q}(\xi_{0},\ldots,\xi_{2m-1})$. Write $\xi_{2m}=\overline{b}/\overline{a}$ with $\overline{a},\overline{b}\in\mathbb{F}_{q}[\xi_{0},\ldots,\xi_{2m-1}]=\operatorname{image}(\Phi_{2m-1,m})$. We may assume that $\overline{a}$ and $\overline{b}$ are relatively prime. Since $\Phi_{2m-1,m}$ is injective, there are unique $a,b\in R_{2m-1}$ with $\Phi_{2m-1,m}(a)=\overline{a}$ and $\Phi_{2m-1,m}(b)=\overline{b}$. Evaluating $\Phi_{2m,m}(aT_{2m}-b)$ gives $\overline{a}\xi_{2m}-\overline{b}=0$. Therefore $aT_{2m}-b$ lies in the kernel of $\Phi_{2m,m}$. For any $F\in R_{2m}$, for $\ell$ sufficiently large we can divide by $aT_{2m}-b$ to get $a^{\ell}F=(aT_{2m}-b)f+r$ with $f,r\in R_{2m-1}$. Applying $\Phi_{2m,m}$ gives
$\Phi_{2m,m}(a^{\ell}F)=\overline{a}^{\ell}\Phi_{2m,m}(F)=\Phi_{2m-1,m}(r).$
Since $\Phi_{2m-1,m}$ is injective we see that $a^{\ell}F$ is in $\ker(\Phi_{2m,m})$ if and only if $a^{\ell}F=(aT_{2m}-b)f$. Since $\overline{a}$ and $\overline{b}$ are relatively prime, $aT_{2m}-b$ is prime in $R_{2m}$. Hence $F\in\ker(\Phi_{2m,m})$ if and only if $F$ is divisible by $aT_{2m}-b$. ∎
Let $\overline{\Phi}_{i,m}$ denote the composition of $\Phi_{i,m}$ followed by projection onto $S_{m}/x_{1}S_{m}$. Let $\sigma$ denote the algebra monomorphism from $S_{m-1}$ to $S_{m}$ which takes $y_{j}$ to $y_{j+1}$ and $x_{j}$ to $x_{j+1}$ for $j=1,\ldots,m-1$. Let $\psi_{1}$ denote the algebra homomorphism from $S_{m}$ to $S_{m}$ which takes $w$ to $w^{q}-wx_{1}^{q-1}$ for $w$ homogeneous of degree one.
@@
(a) $\ker(\psi_{1})=x_{1}S_{m}$ and $\psi_{1}\circ\psi_{1}(f)=\psi_{1}(f)^{q}$.
(b) $\ker(\overline{\Phi}_{i,m})=\ker(\Phi_{i,m-1})=\ker(\psi_{1}\circ\Phi_{i,m}).$
@@
(a) Clearly $x_{1}S_{m}\subseteq\ker(\psi_{1})$. To see that $x_{1}S_{m}=\ker(\psi_{1})$, identify $S_{m}/x_{1}S_{m}$ with $\mathbb{F}_{q}[y_{1},y_{2},\ldots,x_{2}]$ and observe that $\psi_{1}$ maps this ring injectively to $\mathbb{F}_{q}[y_{1}^{q}-y_{1}x_{1}^{q-1},y_{2}^{q}-y_{2}x_{1}^{q-1},\ldots,x_{2}^{q}-x_{2}x_{1}^{q-1}]$. For $w\in S_{m}$ with $\deg(w)=1$, we have $\psi_{1}\circ\psi_{1}(w)=\psi_{1}(w^{q}-wx_{1}^{q-1})=\psi_{1}(w)^{q}$. The general result follows from the fact that taking the $q^{th}$ power is an algebra monomorphism on $S_{m}$.
@@
(b) If we identify $S_{m}/x_{1}S_{m}$ with $\sigma(S_{m-1})[y_{1}]$, then $\overline{\Phi}_{i,m}=\sigma\circ\Phi_{i,m-1}$. Since $\sigma$ is a monomorphism, we see that $\ker(\overline{\Phi}_{i,m})=\ker(\Phi_{i,m-1})$. From (a) we have $\ker(\psi_{1})=x_{1}S_{m}$. Therefore $\ker(\overline{\Phi}_{i,m})=\ker(\psi_{1}\circ\Phi_{i,m})$ follows from the definition of $\overline{\Phi}_{i,m}$.
@@
(c) If $f\in\ker(\overline{\Phi}_{i,m})$ then $x_{1}$ divides $\Phi_{i,m}(f)$. Since $\Phi_{i,m}(f)$ is invariant, if $x_{1}$ divides $\Phi_{i,m}(f)$, each element in the orbit of $x_{1}$ must divide $\Phi_{i,m}(f)$. The result follows from the fact that $u_{m}$ is a product of independent linear factors from the orbit of $x_{1}$

### v3

@@ 3.3. Steenrod Operations
The Steenrod operations are cohomology operations widely used in algebraic topology. They can be sensibly defined on $S_{m}$ and restrict to operations on $S_{m}^{G_{m}}$, see, for example, [24]. Thus applying a Steenrod operation to an invariant produces a potentially new invariant. The complete Steenrod operator $\mathcal{P}(t):S_{m}\to S_{m}[t]$ is the algebra homomorphism determined by $\mathcal{P}(t)(v)=v+v^{q}t$ for $v$ homogeneous of degree one. Since the map is linear in degree one, $\mathcal{P}(t)$ is well-defined. For $f$ homogeneous of degree $d$, the Steenrod operations $\mathcal{P}^{i}(f)$ are defined by
@@
Suppose $v,f\in S_{m}$ with $v$ homogeneous of degree one. Then $v$ divides $\mathcal{P}^{i}(vf)$ for all $i$. Moreover, if $f$ is a product of pairwise relatively prime linear forms, then $f$ divides $\mathcal{P}^{i}(f)$.
@@
By definition, $v$ divides $\mathcal{P}^{j}(v)$. Therefore, using the Cartan identity, $v$ divides $\mathcal{P}^{i}(vf)$. The second assertion follows. ∎
@@
(a) $\mathcal{P}(t)(\xi_{1})=\xi_{1}+\xi_{2}t+\xi_{1}^{q}t^{2}$.
(b) $\mathcal{P}(t)(\xi_{2})=\xi_{2}+2\xi_{1}^{q}t+\xi_{3}t^{q}+\xi_{2}^{q}t^{q+1}$.
(c) For $i\geq 3$, $\mathcal{P}(t)(\xi_{i})=\xi_{i}+\xi_{i-1}^{q}t+\xi_{i+1}t^{q^{i-1}}+\xi_{i}^{q}t^{q^{i-1}+1}$.
@@
(b) $\mathcal{P}^{j+1}(\xi_{1}b)=\xi_{2}b^{q}+\xi_{1}^{q}\mathcal{P}^{j-1}(b)$.
(c) $\mathcal{P}^{j+3}(\xi_{1}^{2}b)=2\xi_{1}^{q}\xi_{2}b^{q}+\xi_{1}^{2q}\mathcal{P}^{j-1}(b)$.
(d) $\mathcal{P}^{2q+j-1}(\xi_{1}^{q}b)=\xi_{1}^{q^{2}}\mathcal{P}^{j-1}b$.
@@ 3.8. New Invariants
Let $P_{m}$ denote the intersection of $G_{m}$ with the upper triangular unipotent subgroup of $\text{GL}_{{n}}(\mathbb{F}_{q})$}}. $P_{m}$ is a choice of Sylow $p$-subgroup for $G_{m}$. For $v\in S_{m}$ homogenous of degree one, define $N(v)$ to be the orbit product of $v$ over $P_{m}$. Since $x_{m}\in S_{m}^{P_{m}}$, $N(x_{m})=x_{m}$. We will see in Section 5 that $\deg(N(x_{i}))=q^{m-i}$ and $\deg(N(y_{i}))=q^{m+i-2}$. Define
@@
Observe that the linear factors of $u_{m}$ lie in a single $G_{m}$-orbit. Furthermore, that orbit consists of the non-zero scalar multiples of the linear factors of $u_{m}$ (see the proof of Theorem 3.10).
Since $u_{m}$ is a product of pairwise relatively prime linear forms, Lemma 3.5 implies that $u_{m}$ divides $\mathcal{P}^{j}(u_{m})$ in $S_{m}$. Define $e(i,m):=\sum_{j=1}^{i}q^{n-1-j}$ with $n=2m$ and, for $1\leq i\leq m$, define $d_{i,m}\in S_{m}$ by
@@
For convenience, define $d_{0,m}=1$. Observe that $e(i,m)=q^{n-2}+qe(i-1,m-1)$ and that the degree of $d_{i,m}$ is $q^{n-1}-q^{n-1-i}$.
@@
$\psi(u_{m})/u_{m}$ is the minimal polynomial of $x_{m}$ over the field $\mathcal{F}(S_{m}^{G_{m}})$. In particular, $d_{i,m}\in S_{m}^{G_{m}}$ for $i=1,\ldots,m$.
@@
If we identify $G_{m-1}$ with the pointwise stabiliser of $\{y_{m},x_{m}\}$ in $G_{m}$, then the pointwise stabiliser of $x_{m}$ coincides with $G_{m-1}H$, where $H$ is the associated Hook group (see Section 5.1). From this we see that the size the orbit of $x_{m}$ is
$\frac{|{{\text{O}^{+}_{{n}}(\mathbb{F}_{q})}}|}{|{{\text{O}^{+}_{{n-2}}(\mathbb{F}_{q})}}||H|}=\frac{(q^{m}-1)(q^{2m-2}-1)}{q^{m-1}-1}=(q^{m}-1)(q^{m-1}+1)=(q-1)\deg(u_{m})$
and this orbit consists of the non-zero scalar multiples of the linear factors of $u_{m}$. If $v$ is a linear factor of $u_{m}$, then the roots of $\psi(v)$ are the non-zero scalar multiples of $v$. From this, using the fact that $\psi(u_{m})/u_{m}$ is monic of degree $(q-1)\deg(u_{m})$, we conclude that $\psi(u_{m})/u_{m}=\prod\{t-x_{m}g\mid g\in G_{m}\}\in S_{m}^{G_{m}}[t]$. Therefore, $\psi(u_{m})/u_{m}$ is the minimal polynomial of $x_{m}$ over $\mathcal{F}(S_{m}^{G_{m}})$ and, since the $d_{i,m}$ are coefficients of $\psi(u_{m})/u_{m}$ up to sign, we have $d_{i,m}\in S_{m}^{G_{m}}$. ∎
@@
$\operatorname{LT}(u_{m})=\prod_{j=1}^{m}y_{j}^{q^{m+j-2}}x_{j}^{q^{m-j}}$
and $\operatorname{LT}(d_{1,m})=y_{m}^{q^{n-1}-q^{n-2}}$.
@@
$u_{m}d_{1,m}=\mathcal{P}^{q^{n-2}}\left(N(y_{m})\right)x_{m}\prod_{i=1}^{m-1}N(y_{i})N(x_{i})+F$
where $\deg_{y_{m}}(F)<q^{n-1}$. Therefore, since $\operatorname{LT}(\mathcal{P}^{q^{n-2}}(N(y_{m}))=y_{m}^{q^{n-1}}$, we see that $\operatorname{LT}(u_{m}d_{1,m})=y^{q^{n-1}-q^{n-2}}\operatorname{LT}(u_{m})$. Dividing by $u_{m}$ gives $\operatorname{LT}(d_{1,m})=y_{m}^{q^{n-1}-q^{n-2}}$. ∎
Define $\mathcal{H}:=\{\xi_{1},\ldots,\xi_{m},d_{1,m},\ldots,d_{m,m}\}$.
@@
$\mathcal{V}(\xi_{1},\ldots,\xi_{m})=\bigcup_{g\in{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}g\mathcal{V}(x_{1},x_{2},\ldots,x_{m}).$
For $v\in\mathcal{V}(\xi_{1},\ldots,\xi_{m})$, choose $g\in{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ so that $gv\in\mathcal{V}(x_{1},x_{2},\ldots,x_{m})$. From Theorem 3.10, $d_{i,m}\in S_{m}^{G_{m}}$. Therefore $d_{i,m}(gv)=d_{i,m}(v)$. Hence to show $\mathcal{V}(\mathcal{H})=\{\underline{0}\}$, it is sufficient to show $\mathcal{V}(x_{1},\ldots,x_{m},d_{1,m},\ldots,d_{m,m})=\{\underline{0}\}$. To do this we consider $d_{i,m}$ modulo the ideal $I:=\langle x_{1},x_{2},\ldots,x_{m}\rangle$.
Define $\bar{d}_{i}\in\mathbb{F}_{q}[y_{1},\ldots,y_{m}]$ by $\bar{d}_{i}\equiv_{I}d_{i,m}$. We will show that $\{\bar{d}_{1},\ldots,\bar{d}_{m}\}$ is an homogeneous system of parameters in $\mathbb{F}_{q}[y_{m},y_{m-1},\ldots,y_{1}]$. Define $W_{m}:={\rm Span}_{\mathbb{F}_{q}}\{y_{m},\ldots,y_{1}\}$. Let $d_{i}(W_{m})$ denote the $i^{\text{th}}$ Dickson invariant in the variables $y_{m},\ldots,y_{1}$. For a definition and background material on the Dickson invariants see [29] or [12] or [5, §3.3]. We claim that
$d_{i,m}\equiv_{I}\pm d_{i}(W_{m})^{q^{m-1}}$
for all $i$. From this, it follows that $\{\bar{d}_{1},\ldots,\bar{d}_{m}\}$ is an homogeneous system of parameters for $\mathbb{F}_{q}[y_{m},\ldots,y_{1}]$, as required.
To see that $d_{i,m}\equiv_{I}\pm d_{i}(W_{m})^{q^{m-1}}$, we look at $\psi(u_{m})/u_{m}$ modulo $I$. Using Theorem 3.10, its proof and the action of the Hook group given in Section 5, we have
$\psi(u_{m})/u_{m}=\prod\{t-x_{m}g\mid g\in G_{m}\}\equiv_{I}t^{q^{m}-1}\left(\prod_{w\in W_{m}\setminus\{0\}}(t-w)\right)^{q^{m-1}}.$
On the right-hand side of this equivalence, $d_{i}(W_{m})^{q^{m-1}}$ is the coefficient of
@@
from which we conclude $d_{i,m}\equiv_{I}\pm d_{i}(W_{m})^{q^{m-1}}$, where the parity is determined by $(-1)^{\ell}=(-1)^{i+1}$. ∎
@@
$\operatorname{LM}(d_{i,m})=\operatorname{LM}(d_{i}(W_{m}))^{q^{m-1}}=\prod_{j=m-i+1}^{m}y_{j}^{q^{m+j-1}-q^{m+j-2}}.$
(The lead monomial of $d_{i}(W_{m})$ can be computed using [29, Proposition 1.3(b)]).
@@ 4. Formulating the Main Theorem
Define $R_{i}:=\mathbb{F}_{q}[T_{i},T_{i-1},\ldots,T_{1}]$ and let $\Phi_{i,m}:R_{i}\to S_{m}$ denote the algebra map defined by $\Phi_{i,m}(T_{j})=\xi_{j}$. If we grade $R_{i}$ by assigning $T_{j}$ degree $q^{j-1}+1$, then $\Phi_{i,m}$ preserves degree. Since $\{\xi_{1},\ldots,\xi_{2m}\}$ is algebraically independent in $S_{m}$, $\Phi_{i,m}$ is injective for $i<2m+1$.
@@
The kernel of $\Phi_{2m+1,m}$ is a principal ideal. Furthermore, the generator is of the form $aT_{2m+1}-b$ with $a,b\in R_{2m}\setminus\{0\}$.
@@
The field of fractions of $S_{m}^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}$ is $\mathbb{F}_{q}(\xi_{1},\ldots,\xi_{2m})$. Write $\xi_{2m+1}=\overline{b}/\overline{a}$ with $\overline{a},\overline{b}\in\mathbb{F}_{q}[\xi_{1},\ldots,\xi_{2m}]=\operatorname{image}(\Phi_{2m,m})$. We may assume that $\overline{a}$ and $\overline{b}$ are relatively prime. Since $\Phi_{2m,m}$ is injective, there are unique $a,b\in R_{2m}$ with $\Phi_{2m,m}(a)=\overline{a}$ and $\Phi_{2m,m}(b)=\overline{b}$. Evaluating $\Phi_{2m+1,m}(aT_{2m+1}-b)$ gives $\overline{a}\xi_{2m+1}-\overline{b}=0$. Therefore $aT_{2m+1}-b$ lies in the kernel of $\Phi_{2m+1,m}$. For any $F\in R_{2m+1}$, for $\ell$ sufficiently large we can divide by $aT_{2m+1}-b$ to get $a^{\ell}F=(aT_{2m+1}-b)f+r$ with $f,r\in R_{2m}$. Applying $\Phi_{2m+1,m}$ gives
$\Phi_{2m+1,m}(a^{\ell}F)=\overline{a}^{\ell}\Phi_{2m+1,m}(F)=\Phi_{2m,m}(r).$
Since $\Phi_{2m,m}$ is injective we see that $a^{\ell}F$ is in $\ker(\Phi_{2m+1,m})$ if and only if $a^{\ell}F=(aT_{2m+1}-b)f$. Since $\overline{a}$ and $\overline{b}$ are relatively prime, $aT_{2m+1}-b$ is prime in $R_{2m+1}$. Hence $F\in\ker(\Phi_{2m+1,m})$ if and only if $F$ is divisible by $aT_{2m+1}-b$. ∎
Let $\overline{\Phi}_{i,m}$ denote the composition of $\Phi_{i,m}$ followed by projection onto $S_{m}/x_{1}S_{m}$. Let $\sigma$ denote the inclusion of $S_{m-1}\subset S_{m}$. Recall that $\psi_{[m,1]}$ is the algebra homomorphism from $S_{m}$ to $S_{m}$ which takes $a$ to $a^{q}-ax_{m}^{q-1}$ for $a$ homogeneous of degree one.
@@
(a) $\ker(\psi_{[m,1]})=x_{m}S_{m}$ and $\psi_{[m,1]}\circ\psi_{[m,1]}(f)=\psi_{[m,1]}(f)^{q}$.
(b) $\ker(\overline{\Phi}_{i,m})=\ker(\Phi_{i,m-1})=\ker(\psi_{[m,1]}\circ\Phi_{i,m}).$
@@
(a) Clearly $x_{m}S_{m}\subseteq\ker(\psi_{[m,1]})$. To see that $x_{m}S_{m}=\ker(\psi_{[m,1]})$, identify $S_{m}/x_{m}S_{m}$ with $\mathbb{F}_{q}[y_{m},y_{m-1},\ldots,x_{m-1}]$ and observe that $\psi_{[m,1]}$ maps this ring injectively to $\mathbb{F}_{q}[y_{m}^{q}-y_{m}x_{m}^{q-1},y_{m-1}^{q}-y_{m-1}x_{m}^{q-1},\ldots,x_{m-1}^{q}-x_{m-1}x_{m}^{q-1}]$. For $a\in S_{m}$ with $\deg(a)=1$, we have $\psi_{[m,1]}\circ\psi_{[m,1]}(a)=\psi_{[m,1]}(a^{q}-ax_{m}^{q-1})=\psi_{[m,1]}(a)^{q}$. The general result follows from the fact that taking the $q^{th}$ power is an algebra monomorphism on $S_{m}$.
@@
(b) If we identify $S_{m}/x_{m}S_{m}$ with $\sigma(S_{m-1})[y_{m}]$, then $\overline{\Phi}_{i,m}=\sigma\circ\Phi_{i,m-1}$. Since $\sigma$ is a monomorphism, we see that $\ker(\overline{\Phi}_{i,m})=\ker(\Phi_{i,m-1})$. From (a) we have $\ker(\psi_{[m,1]})=x_{m}S_{m}$. Therefore $\ker(\overline{\Phi}_{i,m})=\ker(\psi_{[m,1]}\circ\Phi_{i,m})$ follows from the definition of $\overline{\Phi}_{i,m}$.
@@
(c) If $f\in\ker(\overline{\Phi}_{i,m})$ then $x_{m}$ divides $\Phi_{i,m}(f)$. Since $\Phi_{i,m}(f)$ is invariant, if $x_{m}$ divides $\Phi_{i,m}(f)$, each element in the orbit of $x_{m}$ must divide $\Phi_{i,m}(f)$. The result follows from the fact that $u_{m}$ is a product of relatively prime linear factors from the orbit of $x_{m}$